The sum of two numbers is 54. The second is five times the first. Find the numbers?
Answer by stanbon(75887)

The sum of two numbers is 54. The second is five times the first. Find the numbers?
-----------------
Equations:
x + y = 54
y = 5x
-------
Substitute for "y" and solve for "x":
x + 5x = 54
6x = 54
x = 9
---
Since y = 5x, y = 5*9 = 45
================================
Cheers,
Stan H.